# Measuring Compliance of Consent Revocation on the Web

### v2

@@ 9 Recommendations
In this section we propose recommendations to regulators for enhancing the implementation of consent revocation and ensuring legal compliance, in the light of our findings.
@@
R1. Need to unify EU requirements for consent revocation interface. Our research shows that websites implement revocation inconsistently across the Web (see Table IV). For example, the text placed in the website footer differs across websites (e.g., “Cookie Settings”, “Privacy Settings”, but also “EU Privacy”), and in many cases does not mention revocation. It is even harder to locate revocation links within privacy policies that are labelled with text such as “How do I control cookies and how my data is used?”, or “Managing our analytics cookies”. The vague and ambiguous representation of these options misleads the users and prevents them from exercising the revocation right which should be clearly and distinctly recognisable (see LR1). Icons, that are supported by regulators and are easy to locate, use different visualisations, which may also confuse the user. EU DPAs should propose unified interface requirements describing the interface, location, and wordings of revocation. An inspiration can be taken from CPRA regulation in California, where three acceptable wordings are proposed, which simplifies locating it with automated means to measure compliance [41].
@@
R2. Need to standardize consent storage. We observed various inconsistencies in storing the TCStrings and OneTrust consent strings (see Section 7). This is partly due to the usage of different storage options that are not updated simultaneously when consent is revoked. We invite regulators to establish standards for consent storage, and implement security measures to protect the integrity of consent strings.
@@
R3. Need to standardize consent communication via event listeners. Our results show that only 43 out of 163 websites had third-parties installing event listeners in order to be updated about the change in consent decisions (see Section 8.1.1). While OneTrust suggests developers to add event listeners to synchronize consent updates when integrating advertising systems [73], our results show that IAB Europe TCF does not standardize event listeners, prompting every CMP to propose their own solution. This makes it harder for third-parties to adapt to different implementations across websites, placing the burden on them along with legal consequences (for continuous processing of personal data), without offering a practical solution to ensure compliance. We therefore recommend standard-setting bodies like IAB Europe TCF to standardize the implementation of event listeners and the callback functions returned by event listeners to inform third-parties about consent revocation. We also propose regulators to take a position on the means of communicating consent to third-parties. We believe that event listeners are able to provide reasonable means for communicating revocation decisions, and help in the allocation of responsibility to collect valid consent by third-parties present on a website.
@@
R4. Need to regulate consent communication via HTTP requests. The most prevalent violation we detected indicates that 74.2% of websites do not properly inform all third-parties about consent revocation when the HTTP request method is used (see Section 8.1.1). Our result shows that this concrete way of informing third-parties exempts big advertising actors, such as doubleclick.net and criteo.com (see Figure 5) from any responsibility regarding data deletion when users revoke consent. Moreover, these third-parties only receive positive consent from the majority of websites and consequently, while illegally processing data from users who revoked consent, these companies can falsely demonstrate evidence of compliance to regulators by providing a false proof of consent. We therefore propose that EU regulators express their position regarding how websites should inform third-parties, and whether HTTP requests is an acceptable method, since our findings demonstrate that this approach leads to the most prevalent violations.
10 Data Availability & Responsible Disclosure
Data availability. As described in Section 4.1 and Figure 1, we collect background data, screenshots and label the websites during data collection. Screenshots and labels help in categorize the websites into different categories, "via same interface", "via different interface" and "no revocation". A few examples can be found in the Appendix, in Figures 6 and 7. A comprehensive list of these websites is available in the supplementary material containing the dataset and crawler [74]. The network logs to identify inconsistencies and code of the crawler to collect the data are also available in the supplementary material.
Responsible disclosure. As of the date of submission of this paper, 14 November 2024, we have notified all the companies who own the domains that we explicitly mention in the main text of our paper. These domains were mentioned in Sections 57 and 8.1.1. We will add information on the responses we will receive upon acceptance of this paper.
11 Conclusion

### v3

@@ 9. Recommendations
In this section, we also propose recommendations to regulators for enhancing the implementation of consent revocation and ensuring legal compliance, in the light of our findings.
@@
R1. Need to unify EU requirements for consent revocation interface. Our research shows that websites implement revocation inconsistently across the Web (see Table 4). For example, the text placed in the website footer differs across websites (e.g., “Cookie Settings”, “Privacy Settings”, but also “EU Privacy”), and in many cases does not mention revocation. It is even harder to locate revocation links within privacy policies that are labelled with text such as “How do I control cookies and how my data is used?”, or “Managing our analytics cookies”. The vague and ambiguous representation of these options misleads the users and prevents them from exercising the revocation right which should be clearly and distinctly recognisable (see LR1). Icons, that are supported by regulators and are easy to locate, use different visualisations, which may also confuse the user. EU DPAs should propose unified interface requirements describing the interface, location, and wordings of revocation. An inspiration can be taken from CPRA regulation in California, where three acceptable wordings are proposed, which simplifies locating it with automated means to measure compliance (Tran et al., 2024).
@@
R2. Need to standardize consent storage. We observed various inconsistencies in storing the TCStrings and OneTrust consent strings (see §7). This is due to the usage of different storage options that are not updated simultaneously when consent is revoked. We invite regulators to establish standards for consent storage, and implement security measures to protect the integrity of consent strings.
@@
R3. Need to standardize consent communication via event listeners. Our results show that only 43 out of 163 websites had third-parties installing event listeners in order to be updated about the change in consent decisions (see §8.1.1). While OneTrust suggests developers to add event listeners to synchronize consent updates when integrating advertising systems (OneTrust, 2024), our results show that IAB Europe TCF does not standardize event listeners, prompting every CMP to propose their own solution. This makes it harder for third-parties to adapt to different implementations across websites, placing the burden on them along with legal consequences (for continuous processing of personal data), without offering a practical solution to ensure compliance. We therefore recommend standard-setting bodies like IAB Europe TCF to standardize the implementation of event listeners and the callback functions returned by event listeners to inform third-parties about consent revocation. We also propose regulators to take a position on the means of communicating consent to third-parties. We believe that event listeners are able to provide reasonable means for communicating revocation decisions, and help in the allocation of responsibility to collect valid consent by third-parties present on a website.
@@
R4. Need to regulate consent communication via HTTP requests. The most prevalent violation we detected indicates that 74.2% of websites do not properly inform all third-parties about consent revocation when the HTTP request method is used (see §8.1.1). Our result shows that this concrete way of informing third-parties exempts big advertising actors, such as doubleclick.net and criteo.com (see Figure 5) from any responsibility regarding data deletion when users revoke consent. Moreover, these third-parties only receive positive consent from the majority of websites and consequently, while illegally processing data from users who revoked consent, these companies can falsely demonstrate evidence of compliance to regulators by providing a false proof of consent. We therefore propose that EU regulators express their position regarding how websites should inform third-parties, and whether HTTP requests is an acceptable method, since our findings demonstrate that this approach leads to the most prevalent violations.
10. Conclusion
@@
Acknowledgements
This work has been supported by the ANR 22-PECY-0002 IPoP (Interdisciplinary Project on Privacy) project of the Cybersecurity PEPR, the TULIP project of the ANR MRSEI program 2023, and the Inria International Chair funding.
Appendix
Appendix A Inconsistency between consent returned via __tcfapi and consent shared on the network
Responsible party Website Reason for inconsistency
3rd-party (CMP) Ketch Kloud, Inc. forbes.com Different TCString: number of vendor consents and number of vendor legitimate purposes increased in TCString on the network w.r.t to the TCString returned by the __tcfapi.
3rd-party (CMP) Ketch Kloud, Inc. time.com Different TCString: number of vendor consents, number of vendor legitimate purposes and number purpose consent increased in TCString on the network w.r.t TCString returned by the __tcfapi. 13 different TCStrings received in the responses.
1st-party n-tv.de Hardcoded TCString: String hardcoded in 1st-party script, sent in the request to 3rd-party. The TCString had additional purpose legitimate interests (5) and vendor legitimate interests (3).
3rd-party (CMP) Didomi cadenaser.com Third-party sends a different TCString w.r.t TCString returned by TCString in the response to a request made by a TP which was informed of the correct (TCString returned by __tcfapi) in previous requests.
1st-party deadline.com Old, non-updated (after acceptance) consent is shared via network requests even after revocation.
3rd-party (kinja-static.com) kotaku.com Third-party script was loaded after the consent was updated (from the previous stages, both in case of after acceptance and revocation), but it still used the wrong consent.
3rd-party (CMP) InMobi PTE Ltd manchester-eveningnews.co.uk, walesonline.co.uk Although all the requests sent an updated consent after revocation, signal-beacon.s-onetag.com/beacon.min.js script continues to send the positive consent (old consent-after acceptance)
Table 8. Inconsistencies in consent strings shared over the network
Purpose Name Repeated from v2.0 Requires Consent
1 Store and/or access information on a device = 1
2 Use limited data to select advertising Looks new but related to no. 2 = “Select basic ads”
3 Create profiles for personalised advertising = 3
4 Use profiles to select personalised advertising = 4
5 Create profiles to personalise content = 5 (✓) Not clear: in principle, this purpose can be legitimized under a legitimate interest, but it would fail the this test.
6 Use profiles to select personalised content = 6 (✓) Not clear: in principle, this purpose can be legitimized under a legitimate interest, but it would fail the legitimate interest test.
7 Measure advertising performance = 7
8 Measure content performance = 8
9 Understand audiences through statistics or combinations of data from different sources New language but looks like the former purpose no. 9, "Apply market research to generate audience insights"
10 Develop and improve services = 10 It is not specific, and so we cannot derive its legal basis.
X but in principle it could rely on LI though it could fail the LI test
11 Use limited data to select content New X
Table 9. IAB TCF purposes in v2.2 in the TCString and the applicable legal basis. The “Requires Consent” column sums up our analysis. The “Repeated from v2.0.” column compares the purposes of versions v2.0 and v2.2. We add parentheses if exceptions occur. You can find the descriptions of all the purposes in the supplementary material (Anonymous, 2025).
Table 8 presents different types of inconsistencies we have observed in our analysis, grouped by the reason for such inconsistency. We observed three types of inconsistencies in the requests and responses: (1) due to delay in updating the consent after user revokes consent; (2) due to introducing/using a different TCString; (3) due to some scripts not being updated about revocation in consent.
Cases where the consent was not updated immediately are the requests sent to third-parties and containing the TCString before user update. We also investigate the time difference between the user update event and when the old consent was still being sent on the network. We additionally observe cases where some URL parameters were updated while some variable in the POST data containing the old consent and taking few seconds to be updated. Examples include cnn.com, idnes.cz,independent.co.uk.
Cases where a completely different TCString was shared on the network, include two websites forbes.com and time.com, where the CMP’s server sends a TCString in the response to a request made by the CMP’s script. This TCString is not the same as the one returned by the __tcfapi and contains additional vendors and purposes. We however don’t observe this TCString being shared to other third-parties. We observe this behavior both after acceptance and after revocation. In case of time.com, we observed 13 different TCStrings. We also observe one case where a completely different TCString with additional legitimate interests in n-tv.de w.r.t the TCString retuned by the __tcfapi after both in case of after acceptance and after revocation. In this case, the TCString was hard-coded in a first-party script and sent to a third-party. In case of cadenaser.com, a third-party server sends a different TCString in response to a third-party script request. However it was not further sent to other third-parties. We also want to point out here that we observe four cases where the ConsentScreen parameter in the TCString is changed. However it does not affect the user’s consent in any way, hence is not relevant for the discussion.
Finally, we also observe cases where the scripts do not use the updated consent. While in deadline.com, all the requests use the old consent, in two websites, manchestereveningnews.co.uk and walesonline.co.uk some scripts send requests with the updated consent while the one script (signal-beacon.s-onetag.com/beacon.min.js) still sends the old consent on the network. Upon further investigation, we found that this script is used by the CMP but is not the one responsible for updating consent when the user takes action to do so. Table 8 also shows that localStorage was not updated while the cookies and __tcfapi was. It’s possible that the script used the value from local storage instead, resulting in an outdated consent being sent. However, there is no evidence to support this.
(a) Persistent icon
(b) Footer options
(c) Banner on the policy page
(d) Options on other pages
Figure 6. Revocation options within the interface on websites
(a) Browser settings or 3rd-party links
(b) Consent options after login
(c) Contact/Email
(d) Paywalls
(e) Revocation mentioned but not possible
(f) Privacy settings without links
Figure 7. Websites with revocation options via different interfaces or no revocation options
Within the Same Interface Via Different Interfaces [HTML]EA9999No Revocation
Banner No. of Websites Icon or Button (0 Steps) Footer Options (1 Step) Banner on Policy (1 step) Options through Policy Page ($\geq$ 2 steps) Browser set., 3rd-party links After Login Contact/ Email Paywalls Option mentioned but doesn’t work No revocation
Consent Banner 54 4 27 3 18 2 0 0 0 0 0
No Option Banner 8 0 0 0 5 2 1 0 0 0 0
No Banner 43 0 2 5 7 17 0 3 0 0 4 [+5]
Total 105 4 29 8 30 21 1 3 0 0 9
Table 10. Prevalence and type of consent revocation options on 105 websites (rank 1-200) without detected CMPs
Within the Same Interface Via Different Interfaces [HTML]EA9999No Revocation
Banner No. of Websites Icon or Button (0 Steps) Footer Options (1 Step) Banner on Policy (1 step) Options through Policy Page ($\geq$ 2 steps) Browser set., 3rd-party links After Login Contact/ Email Paywalls Option mentioned but doesn’t work No revocation
Consent Banner 54 5 36 3 3 6 1 0 0 0 0
No Option Banner 0 0 0 0 5 2 1 0 0 0 0
No Banner 2 0 0 0 0 0 0 0 0 0 0
Total 56 5 38 3 3 6 1 0 0 0 0
Table 11. Prevalence and type of consent revocation options on 56 websites (rank 1-200) with detected CMPs
Within the Same Interface Via Different Interfaces [HTML]EA9999No Revocation
Banner No. of Websites Icon or Button (0 Steps) Footer Options (1 Step) Banner on Policy (1 step) Options through Policy Page ($\geq$ 2 steps) Browser set., 3rd-party links After Login Contact/ Email Paywalls Option mentioned but doesn’t work No revocation
Consent Banner 214 33 125 0 26 9 0 3 11 3 7
No Option Banner 0 0 0 0 0 0 0 0 0 0 0
No Banner 11 0 8 0 0 1 0 0 0 0 2
Total 225 33 133 0 26 10 0 3 11 3 9
Table 12. Prevalence and type of consent revocation options on 200 to 5000 ranked websites with detected CMPs